In 175 words, explain why, in any period, a country’s net capital inflows equal its trade deficit? Include examples.
Answer #1

Let us first see the equations :

Quantity demanded in forex market = Exports + Capital inflows

Quantity supplied in forex market = Imports + Capital outflows

At equilibrium :

Quantity demanded = Quantity supplied

Exports + Capital inflows = Imports + Capital outflows

So , Exports - Imports + Capital inflows - Capital outflows = 0

(Exports - Imports) + (Capital inflows - Capital outflows) = 0

(CA) + (FA) = 0

Trade deficit = (Exports - Imports) = - A ( say ) ( When imports are greater than export )

Net capital inflow = Capital inflow - outflow = + A ( When inflow is greater than outflow )

- A + ( + A ) = 0 ( Proven )

Capital inflows mean money is coming in, via investment from foreign countries . On the other hand , current account deficits (trade deficits) mean money is going out to foreign countries . Because the sources of cash must equal the uses of cash in balance of payments account , the money going out (trade deficit) will mean there is money coming in some through other way ( foreign investments ) . So country's net capital inflows equal its trade deficit which balances the payments account .